OpenStudy (anonymous):

In a small bag of M&Ms, there are 9 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(you select first an orange candy and then a blue candy). Round your answer to the nearest hundredth.

OpenStudy (anonymous):

would the answer be .23?

OpenStudy (texaschic101):

9/13 * 4/12 = 9/13 * 1/3 = 9/39 = 3/13 = 0.23 You will need to get a second opinion on this because I am not sure if it is correct

OpenStudy (texaschic101):

@Pulsified ....did you do it the same way I did it ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

answer is correct